# WattsOnAI: Measuring, Analyzing, and Visualizing Energy and Carbon Footprint of AI Workloads

### v1

@@ Abstract.
The rapid advancement of AI, particularly large language models (LLMs), has raised significant concerns about the energy use and carbon emissions associated with model training and inference. However, existing tools for measuring and reporting such impacts are often fragmented, lacking systematic metric integration and offering limited support for correlation analysis among them. This paper presents WattsOnAI, a comprehensive software toolkit for the measurement, analysis, and visualization of energy use, power draw, hardware performance, and carbon emissions across AI workloads. By seamlessly integrating with existing AI frameworks, WattsOnAI offers standardized reports and exports fine-grained time-series data to support benchmarking and reproducibility in a lightweight manner. It further enables in-depth correlation analysis between hardware metrics and model performance and thus facilitates bottleneck identification and performance enhancement. By addressing critical limitations in existing tools, WattsOnAI encourages the research community to weigh environmental impact alongside raw performance of AI workloads and advances the shift toward more sustainable ”Green AI” practices. The code is available at https://github.com/SusCom-Lab/WattsOnAI.
copyright: nonedoi: isbn: conference: ; 2025journalyear: ;price:
@@ 1. Introduction
WattsOnAI (Proposed) Full Full Full Full Full Full Full Full Low
@@
By bridging existing research gaps and addressing corresponding challenges, we introduce WattsOnAI and make the following major contributions.
  • Comprehensive Measurement: WattsOnAI unifies real-time monitoring of energy, power, and hardware metrics with carbon emission estimation, offering synchronized tracking throughout AI workload execution. Upon task completion, WattsOnAI generates intuitive post-task reports and exports detailed time-series data, addressing key gaps in metric integration and carbon-aware reporting found in existing tools.
  • Ease of Use and Low Overhead: WattsOnAI is simple to use and seamlessly integrates with various AI workflows; concurrently, it operates with minimal performance overhead (Section 4). This addresses the challenge of a high barrier to entry, avoids impacting AI task performance, and ensures reliable long-term measurement.
  • Multi-Dimensional Visualization and Correlation Analysis: WattsOnAI implements multi-dimensional visualization capabilities, effectively demonstrating how metrics such as energy consumption, carbon emissions, and hardware characteristics evolve over time and relate to one another. Through these tailored visualizations, correlation analysis across classified metrics spanning compute, memory, and communication dimensions is enabled. This bridges a key gap in existing tools by supporting multi-dimensional, workload-aware efficiency analysis.
@@ 2.1. Hardware Monitoring Interfaces and Libraries
Figure 1. The architecture of WattsOnAI.
@@ 3. Architecture and Design
Fig. 1 shows the architecture of WattsOnAI, which is fundamentally divided into i) source interface layer, ii) data processing and persistance layer, and iii) demonstration layer. Through this three-layer architecture, WattsOnAI systematically measures, analyses and visualizes energy consumption, power consumption, carbon emissions and other relevant hardware metrics throughout the execution of AI tasks.
@@ 3.1. Source Interface Layer
As mentioned in the previous section, metric data collection primarily relies on interfaces provided by the hardware. Based on user configuration, WattsOnAI dynamically selects the metrics to sample and can support up to 26 distinct metrics. These include GPU power consumption, SM active, among others.
@@
Given that different hardware components expose metrics through distinct interfaces, and even a single component like an NVIDIA GPU might offer data via multiple mechanisms (e.g., NVML/SMI and DCGM), we employ a parallelized collection strategy. This is handled within WattsOnAI by our MetricsCollector class, which is responsible for the statistical metric collection. Where appropriate (primarily for GPU metrics), this allows WattsOnAI to query interfaces like those underlying SMI and DCGM concurrently. This approach maximizes the achievable sampling rate to approximately 0.1s-0.2s per sample, a frequency no lower than that of pynvml (pynvml, ).
@@ 3.2. Data Processing and Persistence Layer
While embodied carbon is crucial for comparing different hardware solutions or for full Life Cycle Assessments (LCA), WattsOnAI’s primary focus is on quantifying the direct impact of specific AI tasks (training, inference, etc.) during their execution. Consequently, we deliberately exclude embodied carbon and measure only the operational carbon generated during task execution, using the following rule:
@@
The second consideration is the type of carbon intensity factor used. Average intensity reflects the grid’s overall generation mix and is suited for broad estimations. Marginal intensity, however, represents the emissions from the power source(s) activated to meet the additional load imposed by the task. Since this marginal generation often comes from more carbon-intensive sources (like fossil fuel peaker plants), its intensity is typically higher and more accurately reflects the direct environmental consequence of running the AI task. To provide a clearer picture of the incremental impact and environmental cost, WattsOnAI uses marginal carbon intensity for its estimations, with the necessary intensity factors being retrieved from services like Electricity Maps (electricity_maps, ) and WattTime (Watttime, ) based on geographical coordinates.
@@ 3.3. Demonstration Layer
  • Real-time Monitoring Dashboard: We also integrate WattsOnAI with visualization tools (e.g., Grafana (Grafana, )) for live monitoring during task execution.
@@ 4. Case Study
from WattsOnAI import monitor
@@
Utilizing WattsOnAI as demonstrated in the code snippet above, we conducted an LLM inference experiment using the Llama2-7b model on a server with NVIDIA A800 GPU, and sampled All Metrics (see Appendix A for the full list) at a sampling interval of $0.1$ second. With the help of WattsOnAI, this experiment yields findings with the following insights and implications.
@@ 4.2. Analysis on Carbon Emission Estimation
In the aforementioned experiment, the report generated by WattsOnAI indicated that the task consumed 7485 Joules (2.08 $Watts\cdot{hour}$) of energy. However, as we have mentioned, such energy consumption figure does not fully reflect the task’s actual environmental impact, which motivated our introduction of the carbon emissions metric.
@@
4.3. Overhead Caused by WattsOnAI
We evaluated the system overhead of WattsOnAI using pidstat across two control groups. The first group measured the inherent overhead of WattsOnAI by running it on an otherwise idle system with a 0.1-second sampling interval. The second group replicated the experiment from Section 4, but with the modification that we removed the 15-second sleep time between the prefill and decode phases. In this setup, we evaluated the impact of WattsOnAI with sampling intervals of 0.1, 0.5, 1, and 5 seconds. The results are presented in Table 2.
Table 2. Overhead Caused by WattsOnAI (WOA)
@@ 4.3. Overhead Caused by WattsOnAI
WOA(0.1) only 15.05s 0.33%  $\uparrow$ 0.37ms 2.23% 121.29MB
Exp w/o WOA 29.15s 99.1% 1411.37MB
Exp w WOA(5) 29.39s 0.82%  $\uparrow$ 38.51ms 97.21% 1388.59MB
Exp w WOA(1) 29.53s 1.30%  $\uparrow$ 12.46ms 97.67% 1398.04MB
Exp w WOA(0.5) 29.71s 1.92%  $\uparrow$ 9.28ms 98.24% 1398.84MB
Exp w WOA(0.1) 30.54s 4.77%  $\uparrow$ 7.2ms 98.97% 1403.62MB
The data in the table indicates that the overhead caused by WattsOnAI is low. Besides, it can be seen that the higher the sampling frequency, the higher the overhead, but the time consumption per sample is smaller. Regarding CPU utilization, the highest usage is observed when WattsOnAI is absent, as the tool itself consumes a small portion of CPU resources. Finally, the differences in memory consumption are negligible and can be considered measurement noise.
@@ 5. Conclusion and future work
Aiming to foster greater awareness of energy consumption in AI research, we developed a software toolkit WattsOnAI, which enables measuring, analyzing, and visualizing energy and carbon footprint of AI workloads. WattsOnAI also forms a solid foundation for AI energy predicting models. Future work will focus on achieving task-level granularity (e.g., attributing energy consumption to individual processes running concurrently on shared hardware) and extending compatibility beyond NVIDIA GPUs.

### v2

# AIMeter: Measuring, Analyzing, and Visualizing Energy and Carbon Footprint of AI Workloads

@@ Abstract.
The rapid advancement of AI, particularly large language models (LLMs), has raised significant concerns about the energy use and carbon emissions associated with model training and inference. However, existing tools for measuring and reporting such impacts are often fragmented, lacking systematic metric integration and offering limited support for correlation analysis among them. This paper presents AIMeter, a comprehensive software toolkit for the measurement, analysis, and visualization of energy use, power draw, hardware performance, and carbon emissions across AI workloads. By seamlessly integrating with existing AI frameworks, AIMeter offers standardized reports and exports fine-grained time-series data to support benchmarking and reproducibility in a lightweight manner. It further enables in-depth correlation analysis between hardware metrics and model performance and thus facilitates bottleneck identification and performance enhancement. By addressing critical limitations in existing tools, AIMeter encourages the research community to weigh environmental impact alongside raw performance of AI workloads and advances the shift toward more sustainable ”Green AI” practices. The code is available at https://github.com/SusCom-Lab/AIMeter.
copyright: noneconference: ; 2025
@@ 1. Introduction
AIMeter (Proposed) Full Full Full Full Full Full Full Full Low
@@
By bridging existing research gaps and addressing corresponding challenges, we introduce AIMeter and make the following major contributions.
  • Comprehensive Measurement: AIMeter unifies real-time monitoring of energy, power, and hardware metrics with carbon emission estimation, offering synchronized tracking throughout AI workload execution. Upon task completion, AIMeter generates intuitive post-task reports and exports detailed time-series data, addressing key gaps in metric integration and carbon-aware reporting found in existing tools.
  • Ease of Use and Low Overhead: AIMeter is simple to use and seamlessly integrates with various AI workflows; concurrently, it operates with minimal performance overhead (Section 4). This addresses the challenge of a high barrier to entry, avoids impacting AI task performance, and ensures reliable long-term measurement.
  • Multi-Dimensional Visualization and Correlation Analysis: AIMeter implements multi-dimensional visualization capabilities, effectively demonstrating how metrics such as energy consumption, carbon emissions, and hardware characteristics evolve over time and relate to one another. Through these tailored visualizations, correlation analysis across classified metrics spanning compute, memory, and communication dimensions is enabled. This bridges a key gap in existing tools by supporting multi-dimensional, workload-aware efficiency analysis.
@@ 2.1. Hardware Monitoring Interfaces and Libraries
Figure 1. The architecture of AIMeter.
@@ 3. Architecture and Design
Fig. 1 shows the architecture of AIMeter, which is fundamentally divided into i) source interface layer, ii) data processing and persistence layer, and iii) demonstration layer. Through this three-layer architecture, AIMeter systematically measures, analyzes and visualizes energy consumption, power consumption, carbon emissions and other relevant hardware metrics throughout the execution of AI tasks.
@@ 3.1. Source Interface Layer
As mentioned in the previous section, metric data collection primarily relies on interfaces provided by the hardware. Based on user configuration, AIMeter dynamically selects the metrics to sample and can support up to 26 distinct metrics. These include GPU power consumption, SM active, among others.
@@
Given that different hardware components expose metrics through distinct interfaces, and even a single component like an NVIDIA GPU might offer data via multiple mechanisms (e.g., NVML/SMI and DCGM), we employ a parallelized collection strategy. This is handled within AIMeter by our MetricsCollector class, which is responsible for the statistical metric collection. Where appropriate (primarily for GPU metrics), this allows AIMeter to query interfaces like those underlying SMI and DCGM concurrently. This approach maximizes the achievable sampling rate to approximately 0.1s-0.2s per sample, a frequency no lower than that of pynvml (pynvml, ).
@@ 3.2. Data Processing and Persistence Layer
While embodied carbon is crucial for comparing different hardware solutions or for full Life Cycle Assessments (LCA), AIMeter’s primary focus is on quantifying the direct impact of specific AI tasks (training, inference, etc.) during their execution. Consequently, we deliberately exclude embodied carbon and measure only the operational carbon generated during task execution, using the following rule:
@@
The second consideration is the type of carbon intensity factor used. Average intensity reflects the grid’s overall generation mix and is suited for broad estimations. Marginal intensity, however, represents the emissions from the power source(s) activated to meet the additional load imposed by the task. Since this marginal generation often comes from more carbon-intensive sources (like fossil fuel peaker plants), its intensity is typically higher and more accurately reflects the direct environmental consequence of running the AI task. To provide a clearer picture of the incremental impact and environmental cost, AIMeter uses marginal carbon intensity for its estimations, with the necessary intensity factors being retrieved from services like Electricity Maps (electricity_maps, ) and WattTime (Watttime, ) based on geographical coordinates.
@@ 3.3. Demonstration Layer
  • Real-time Monitoring Dashboard: We also integrate AIMeter with visualization tools (e.g., Grafana (Grafana, )) for live monitoring during task execution.
@@ 4. Case Study
from AIMeter import monitor
@@
Utilizing AIMeter as demonstrated in the code snippet above, we conducted an LLM inference experiment using the Llama2-7b model on a server with NVIDIA A800 GPU, and sampled All Metrics (see Appendix A for the full list) at a sampling interval of $0.1$ second. With the help of AIMeter, this experiment yields findings with the following insights and implications.
@@ 4.2. Analysis on Carbon Emission Estimation
In the aforementioned experiment, the report generated by AIMeter indicated that the task consumed 7485 Joules (2.08 $Watts\cdot{hour}$) of energy. However, as we have mentioned, such energy consumption figure does not fully reflect the task’s actual environmental impact, which motivated our introduction of the carbon emissions metric.
@@
4.3. Overhead Caused by AIMeter
We evaluated the system overhead of AIMeter using pidstat across two control groups. The first group measured the inherent overhead of AIMeter by running it on an otherwise idle system with a 0.1-second sampling interval. The second group replicated the experiment from Section 4, but with the modification that we removed the 15-second sleep time between the prefill and decode phases. In this setup, we evaluated the impact of AIMeter with sampling intervals of 0.1, 0.5, 1, and 5 seconds. The results are presented in Table 2.
Table 2. Overhead Caused by AIMeter (AIM)
@@ 4.3. Overhead Caused by AIMeter
AIM(0.1) only 15.05s 0.33%  $\uparrow$ 0.37ms 2.23% 121.29MB
Exp w/o AIM 29.15s 99.1% 1411.37MB
Exp w AIM(5) 29.39s 0.82%  $\uparrow$ 38.51ms 97.21% 1388.59MB
Exp w AIM(1) 29.53s 1.30%  $\uparrow$ 12.46ms 97.67% 1398.04MB
Exp w AIM(0.5) 29.71s 1.92%  $\uparrow$ 9.28ms 98.24% 1398.84MB
Exp w AIM(0.1) 30.54s 4.77%  $\uparrow$ 7.2ms 98.97% 1403.62MB
The data in the table indicates that the overhead caused by AIMeter is low. Besides, it can be seen that the higher the sampling frequency, the higher the overhead, but the time consumption per sample is smaller. Regarding CPU utilization, the highest usage is observed when AIMeter is absent, as the tool itself consumes a small portion of CPU resources. Finally, the differences in memory consumption are negligible and can be considered measurement noise.
@@ 5. Conclusion and future work
Aiming to foster greater awareness of energy consumption in AI research, we developed a software toolkit AIMeter, which enables measuring, analyzing, and visualizing energy and carbon footprint of AI workloads. AIMeter also forms a solid foundation for AI energy predicting models. Future work will focus on achieving task-level granularity (e.g., attributing energy consumption to individual processes running concurrently on shared hardware) and extending compatibility beyond NVIDIA GPUs.